A 34.0-litre cylinder contains 212g of oxygen at 27-degree Celcius.What mass of oxygen must be released to reduce the pressure to 2.463 atm? 

Grade:11

1 Answers

Arun
6 years ago
Deat Ayush
 
moles O2 = 212 g/32 g/mol=6.625

p = 6.625 x 0.08206 x 300 K/ 34.0= 4.7968 = 4.8 atm 


4.8 x 34.0 = 6.625 RT 
2.463 x 34.0 = nRT 

163.2 = 6.625 RT 
83.742 = nRT 

163.2/ 83.742 = 6.625 /n 

n = 3.399 = 3.4

6.625 - 3.4 =3.225 

mass O2 that must be released = 3.225 mol x 32 g/mol=103.2 g
 
Regards,
Arun (AskIITians Forum Expert)